# Range Retrieval with Graph-Based Indices

### v2

@@ Acknowledgements.
- Zhang et al. (2019) Hongfei Zhang, Xia Song, Chenyan Xiong, Corby Rosset, Paul N. Bennett, Nick Craswell, and Saurabh Tiwary. 2019. Generic Intent Representation in Web Search. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2019, Paris, France, July 21-25, 2019, Benjamin Piwowarski, Max Chevalier, Éric Gaussier, Yoelle Maarek, Jian-Yun Nie, and Falk Scholer (Eds.). ACM, 65–74. https://doi.org/10.1145/3331184.3331198
Appendix A Additional Data
We provide data on early stopping metrics for four additional datasets in Figures 10, 11, 12, and 13.
@@ Appendix A Additional Data
In Figure 16 we present further study on size scaling using the Wikipedia dataset at the 1M and 10M scale.

### v3

@@ Acknowledgements.
- Xu et al. (2024) Yuexuan Xu, Jianyang Gao, Yutong Gou, Cheng Long, and Christian S. Jensen. 2024. iRangeGraph: Improvising Range-dedicated Graphs for Range-filtering Nearest Neighbor Search. Proc. ACM Manag. Data 2, 6 (2024), 239:1–239:26. https://doi.org/10.1145/3698814
- Zhang et al. (2019) Hongfei Zhang, Xia Song, Chenyan Xiong, Corby Rosset, Paul N. Bennett, Nick Craswell, and Saurabh Tiwary. 2019. Generic Intent Representation in Web Search. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2019, Paris, France, July 21-25, 2019. ACM, 65–74. https://doi.org/10.1145/3331184.3331198
- Zuo et al. (2024) Chaoji Zuo, Miao Qiao, Wenchao Zhou, Feifei Li, and Dong Deng. 2024. SeRF: Segment Graph for Range-Filtering Approximate Nearest Neighbor Search. Proc. ACM Manag. Data 2, 1 (2024), 69:1–69:26. https://doi.org/10.1145/3639324
Appendix A Supplementary Material
A.1. Additional Data on Early Stopping
We provide data on early stopping metrics for four datasets in Figures 8, 9, 10, and 11. In each figure, we histogram the values of each metric, first with all results included, and then with only points where the search has not already found a point inside the radius. The figure also shows vertical lines indicating a potential cutoff for Algorithm 4; points with more than one range result to the right of the line would be “incorrectly” cut off from further search, while points with zero points to the right of the line would be “incorrectly” allowed to proceed instead of terminating early. The goal is to identify a cutoff that yields a more favorable QPS/recall curve.
@@ A.1. Additional Data on Early Stopping
A.2. Parameter Values
| Dataset | Parameters |
| BIGANN-1M | ”OPQ32_64,IVF65536_HNSW32,PQ64x4fsr” |
| BIGANN-10M | ”OPQ32_64,IVF1048576_HNSW32,PQ64x4fsr” |
| BIGANN-100M | ”OPQ64_128,IVF1048576_HNSW32,PQ128x4fsr” |
| DEEP-1M | ”OPQ32_64,IVF65536_HNSW32,PQ64x4fsr” |
| MSTuring-1M | ”OPQ32_64,IVF16384_HNSW32,PQ64x4fsr” |
| GIST-1M | ”OPQ32_64,IVF16384_HNSW32,PQ64x4fsr” |
| SSNPP-1M | ”OPQ32_64,IVF65536_HNSW32,PQ64x4fsr” |
| SSNPP-10M | ”OPQ64_128,IVF65536_HNSW32,PQ128x4fsr” |
| SSNPP-100M | ”OPQ32_128,IVF1048576_HNSW32,PQ32” |
| OpenAI-1M | ”OPQ32_64,IVF16384_HNSW32,PQ64x4fsr” |
| Text2Image-1M | ”OPQ64_128,IVF16384_HNSW32,PQ128x4fsr” |
| Wikipedia-1M | ”IVF16384_HNSW32,Flat” |
| MSMARCOWebSearch-1M | ”IVF16384_HNSW32,Flat” |
In Figure 14, we show our choice of FAISS parameters for each dataset. For each dataset, we chose a starting set of parameters based on both instructions from the FAISS Wiki (Douze, 2022) and the parameter settings from the NeurIPS’21 Big ANN Benchmarks Competition (Simhadri et al., 2021b). In each case, we use two-level clustering to significantly speed up build times. We then experimented with increasing and decreasing the choice of both the size of the PQ codes and the number of centroids, and kept the parameters that gave the optimal QPS/Average Precision curve. For SSNPP-100M, we also made use of the Hamming threshold parameter as suggested from the configurations in the Big ANN Benchmarks repository. The curves were drawn by varying the HNSW ef_search parameter as well as the number of centroids probed. FAISS has a build-in range search method (Douze, 2022), which we used to conduct searches. For Wikipedia and MSMARCOWebSearch, we used no quantization during range search as the quantization seemed to distort distances for inner product datasets and produced nonsensical results.
| Dataset | Build Parameters | Early Stopping Threshold |
| BIGANN (all sizes) | 10000 | |
| DEEP-1M | .4 | |
| MSTuring-1M | 1.6 | |
| GIST-1M | 2.5 | |
| SSNPP (all sizes) | 200000 | |
| OpenAI-1M | .28 | |
| Text2Image-1M | -1.0 | |
| Wikipedia-1M | -9.5 | |
| MSMARCOWebSearch-1M | -57 |
In Figure 15 we show the choices of build and search parameters for the DiskANN graphs used in our experiments. For the build parameters, we used wherever possible, although we found it was not sufficient for SSNPP. The parameter was configured experimentally for each dataset, as was the choice of early stopping threshold.